# Pascal-Weighted Genetic Algorithms: A Binomially-Structured Recombination Framework

### v1

@@ Abstract
This paper introduces a new family of multi-parent recombination operators for Genetic Algorithms (GAs), based on normalized Pascal (binomial) coefficients. Unlike classical two-parent crossover operators, Pascal-Weighted Recombination (PWR) forms offsprings as structured convex combination of multiple parents, using binomially shaped weights that emphasize central inheritance while suppressing disruptive variance. We develop a mathematical framework for PWR, derive variance-transfer properties, and analyze its effect on schema survival. The operator is extended to real-valued, binary/logit, and permutation representations.
[…]
@@ 1 Introduction
A key observation motivating this work is that the distribution of parental influence should neither be uniform nor dominated by a single parent. Among the many possible weighting strategies, rows of Pascal’s triangle offer a principled middle ground: a smooth, symmetric, unimodal distribution that favors central parents without suppressing peripheral ones. The corresponding binomial weights are mathematically grounded, easy to compute, and naturally scalable with the number of parents, yielding convex offsprings that inherit traits through a balanced and continuous interpolation of parental contributions. Furthermore, Pascal-derived weighting is also more biologically faithful. Whereas standard GA operators emulate a strict two-parent reproductive model, biological heredity is cumulative: traits arise from convex mixtures of influences aggregated across multiple generations [30]. Binomial attenuation, encoded by Pascal coefficients, mirrors this multi-generational inheritance pattern by producing polynomially smooth and variance-controlled recombination surfaces.
@@ 2 Research Gap and Contributions
Although multi-parent recombination (MPR) in genetic algorithms was explored in the 1990s, notably by Eiben et al. [6], and later in studies on real-coded genetic algorithms and advanced crossover schemes [3, 4, 24, 25], the operator design space has remained largely underexplored in the context of modern engineering optimization. Early work focused primarily on heuristic extensions of two-parent crossover, without establishing a principled framework for weight allocation, offspring distribution, or statistical control of diversity and convergence [39].
@@ 4.1 Variance and Diversity Properties
For $m>2$, this quantity is strictly less than $\sigma_{p}^{2}$ and also less than the variance attained by equal-weight averaging. Thus, PWR inherently produces a variance contraction that strengthens with increasing $m$, providing a natural self-regularizing mechanism without additional hyperparameters, analogous in spirit to variance-minimizing information fusion in estimation and filtering [29, 31].
@@
Kalman-like interpretation
Interestingly, the behavior of PWR resembles the variance–minimizing structure of the Kalman filter. In Kalman estimation, multiple uncertain measurements are fused through convex weights chosen to minimize posterior variance. PWR exhibits an analogous mechanism: the binomial weights emphasize central parents (high information content) and down-weight peripheral ones (higher uncertainty), producing an offspring that represents a low-variance fusion of parental information. This provides an intuitive statistical justification for the stability and smooth convergence observed in the experiments [31].
@@ 5.1 Connection to Bernstein Polynomials
$[cc]B(1/2)$ $=$ $\sum_{i=0}^{m-1}\binom{m-1}{i}\left(\tfrac{1}{2}\right)^{i}\left(1-\tfrac{1}{2}\right)^{m-1-i}P_{i+1}$ (4)
@@ Results:
As depicted in Table 1, PWR-3 achieves the lowest ITAE with reduced run-to-run variance. Step responses exhibit faster rise time, smaller overshoot, and smoother settling, confirming that variance-controlled recombination is advantageous for continuous control design [32].
@@ 6.3.2 Traveling Salesman Problem (TSP)
As depected in Table 4, PWR-3 reduces variance and yields shorter average tour lengths while avoiding the disruptive swaps of PMX, suggesting that structured multi-parent influence can be advantageous in permutation spaces [12, 34, 13, 14].
@@ 7.2 Weight Shape
We compare Pascal weights, equal weights ($w_{i}=1/m$), and random Dirichlet-distributed weights. Pascal weights produce the lowest variance and the most stable convergence. Equal weights are moderately effective but less stable in sensitive continuous domains, while random weights are inconsistent and prone to catastrophic updates [26, 23].
@@ 8 Discussion
9 Pascal, Fibonacci, and the Golden Ratio
The Pascal-weighted recombination operator is built on binomial coefficients $\binom{n}{k}$, which appear in numerous combinatorial and analytical contexts. Two particularly important structures related to binomial coefficients are the Fibonacci sequence and the golden ratio $\varphi$. Although the present paper does not rely on these connections for the definition of PWR, it is useful to highlight how they provide additional mathematical context and a possible link to deeper hereditary patterns [30].
9.1 Fibonacci Numbers in Pascal’s Triangle
Let $(F_{n})_{n\geq 0}$ denote the Fibonacci sequence,
$F_{0}=0,\quad F_{1}=1,\quad F_{n+1}=F_{n}+F_{n-1}.$
It is well known that Fibonacci numbers can be read directly from Pascal’s triangle by summing entries along the shallow diagonals. In particular,
$F_{n+1}=\sum_{k=0}^{\lfloor n/2\rfloor}\binom{n-k}{k}$, (5)
so that the same binomial coefficients driving Pascal-weighted recombination also generate the Fibonacci sequence through structured aggregation.
Equation (5) shows that Fibonacci growth can be expressed as a particular superposition of binomial coefficients; conversely, each row of Pascal’s triangle can be seen as a building block for Fibonacci-like recurrence. From an algorithmic standpoint, this suggests that the binomial patterns used in PWR are not isolated artifacts but belong to a broader family of combinatorial structures that also give rise to canonical growth sequences in mathematics and biology [30].
9.2 Golden-Ratio Limit of Fibonacci Ratios
A second well-known fact is that the ratio of successive Fibonacci numbers converges to the golden ratio
$\varphi=\frac{1+\sqrt{5}}{2}\approx 1.618,\qquad\lim_{n\to\infty}\frac{F_{n+1}}{F_{n}}=\varphi.$
The golden ratio appears in diverse models of growth, spacing, and efficient packing in biological and physical systems. In the context of Eq. (5), this implies that certain aggregate binomial sums inherit a characteristic scaling behavior governed by $\varphi$.
Although PWR uses normalized Pascal rows rather than Fibonacci weights explicitly, the underlying binomial structure is mathematically adjacent to Fibonacci and golden-ratio phenomena: both arise from simple recurrences and combinatorial aggregation rules. One may interpret Fibonacci-based scaling as a global manifestation of structures that are, at a more local level, binomially composed.
9.3 Implications for Recombination Design
From the perspective of genetic algorithms, the connection between binomial coefficients, Fibonacci recurrence, and the golden ratio has two main implications:
  • Structured aggregation. Fibonacci numbers can be constructed from binomial coefficients via Eq. (5). This reinforces the view that Pascal-weighted recombination is not an arbitrary choice of weights but a canonical way to aggregate discrete contributions through a simple, universal combinatorial rule.
  • Potential for ratio-based extensions. Since $\varphi$ emerges as the asymptotic ratio of Fibonacci numbers, one can envision golden-ratio-inspired variants of PWR in which the number of parents, or the relative emphasis between central and peripheral parents, is scheduled according to Fibonacci or golden-ratio progressions. Such designs could provide a principled way to modulate the trade-off between exploration and exploitation across generations.
These connections are primarily conceptual in the present work: PWR, as evaluated in this paper, relies only on normalized Pascal rows. However, the shared combinatorial foundation with Fibonacci and golden-ratio structures suggests a broader research direction in which recombination operators are derived not only from binomial coefficients but also from other mathematically structured sequences with known scaling and stability properties.

### v2

@@ Abstract
This paper introduces a new family of multi-parent recombination operators for Genetic Algorithms (GAs), based on normalized Pascal (binomial) coefficients. Unlike classical two-parent crossover operators, Pascal-Weighted Recombination (PWR) forms offspring as structured convex combination of multiple parents, using binomially shaped weights that emphasize central inheritance while suppressing disruptive variance. We develop a mathematical framework for PWR, derive variance-transfer properties, and analyze its effect on schema survival. The operator is extended to real-valued, binary/logit, and permutation representations.
We evaluate the proposed method on four representative benchmarks: (i) PID controller tuning evaluated using the ITAE metric, (ii) FIR low-pass filter design under magnitude-response constraints, (iii) wireless power-modulation optimization under SINR coupling, and (iv) the Traveling Salesman Problem (TSP). We demonstrate how, across these benchmarks, PWR consistently yields smoother convergence, reduced variance, and often achieves measurable performance improvements (up to 9–22% in selected benchmarks) with reduced run-to-run variance. The approach is simple, algorithm-agnostic, and readily integrable into diverse GA architectures.
@@ 1 Introduction
A key observation motivating this work is that the distribution of parental influence should neither be uniform nor dominated by a single parent. Among the many possible weighting strategies, rows of Pascal’s triangle offer a principled middle ground: a smooth, symmetric, unimodal distribution that favors central parents without suppressing peripheral ones. The corresponding binomial weights are mathematically grounded, easy to compute, and naturally scalable with the number of parents, yielding convex offspring that inherit traits through a balanced and continuous interpolation of parental contributions. Furthermore, Pascal-derived weighting is also more biologically faithful. Whereas standard GA operators emulate a strict two-parent reproductive model, biological heredity is cumulative: traits arise from convex mixtures of influences aggregated across multiple generations [30]. Binomial attenuation, encoded by Pascal coefficients, mirrors this multi-generational inheritance pattern by producing polynomially smooth and variance-controlled recombination surfaces.
@@ 2 Research Gap and Contributions
Although multi-parent recombination (MPR) in genetic algorithms was explored in the 1990s, notably by Eiben et al. [6], and later in studies on real-coded genetic algorithms and advanced crossover schemes [3, 24, 25], the operator design space has remained largely underexplored in the context of modern engineering optimization. Early work focused primarily on heuristic extensions of two-parent crossover, without establishing a principled framework for weight allocation, offspring distribution, or statistical control of diversity and convergence [39].
@@ 4.1 Variance and Diversity Properties
For $m>2$, this quantity is strictly less than $\sigma_{p}^{2}$, so Pascal-weighted mixing contracts variance relative to selecting a single parent.
It is important to note, however, that among all convex combinations with $\sum_{i}w_{i}=1$, the minimum of $\sum_{i}w_{i}^{2}$ (and hence the minimum offspring variance in Eq. (3)) is achieved by equal weights $w_{i}=1/m$. Pascal weights therefore do not minimize variance; rather, they impose a smooth, symmetric central-emphasis profile (interpretable via Bernstein/Bézier structure) that trades a small amount of additional variance for structured inheritance.
@@ 5.1 Connection to Bernstein Polynomials
$B(1/2)$ $=\sum_{i=0}^{m-1}\binom{m-1}{i}\left(\tfrac{1}{2}\right)^{i}\left(1-\tfrac{1}{2}\right)^{m-1-i}P_{i+1}$
@@ Results:
As depicted in Table LABEL:table1, PWR-3 achieves the lowest ITAE with reduced run-to-run variance. Step responses exhibit faster rise time, smaller overshoot, and smoother settling, confirming that variance-controlled recombination is advantageous for continuous control design [32].
@@ 6.3.2 Traveling Salesman Problem (TSP)
As depicted in Table 4, PWR-3 reduces variance and yields shorter average tour lengths while avoiding the disruptive swaps of PMX, suggesting that structured multi-parent influence can be advantageous in permutation spaces [12, 34, 13, 14].
@@ 7.2 Weight Shape
We compare Pascal weights, equal weights ($w_{i}=1/m$), and random Dirichlet-distributed weights. Equal weights minimize $\sum_{i}w_{i}^{2}$ among convex combinations and therefore yield the lowest offspring variance in Eq. (3), but they can dilute fitness information by treating all parents symmetrically. Pascal weights impose a smooth, symmetric central-emphasis profile (interpretable via Bernstein/Bézier structure) that contracts variance relative to single-parent inheritance while preserving structured inheritance from central parents, leading to more stable convergence in practice. Random Dirichlet-distributed weights exhibit high run-to-run variability and can produce disruptive updates [23, 26].